OpenStudy (anonymous):

Show that:

OpenStudy (anonymous):

\[\frac{ \log_mx }{ \log _{mn}x}=1+\log_mn\]

OpenStudy (anonymous):

@ganeshie8 @Miracrown

ganeshie8 (ganeshie8):

Take the left hand side and use change of base formula

OpenStudy (anonymous):

what does it become?

OpenStudy (anonymous):

do you know?

ganeshie8 (ganeshie8):

\[\log_ba = \dfrac{\ln ~a}{\ln~ b}\]

OpenStudy (anonymous):

log m xlog mn x =1+log m n

OpenStudy (anonymous):

I don't understand what to do with the base mn

OpenStudy (anonymous):

@ganeshie8

OpenStudy (triciaal):